# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ D.2 Proofs of Lemmas for Theorem 1
$\leq\frac{4K^{2}}{\kappa}d\log\left(1+\frac{2KT}{d\lambda}\right).$ (Lemma D.3)
On the other hand, for $t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}={\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$, we know that
$\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}\geq\frac{|\widetilde{{\mathcal{T}}}^{w}_{T+1}|}{3\sqrt{2}\beta_{T+1}(\delta)^{2}}.$
By combining the two results above, we obtain
@@
$|\widetilde{{\mathcal{T}}}^{w}_{T+1}|=|{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}|\leq\frac{12\sqrt{2}K^{2}}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right),$
@@
Appendix E Proof of Theorem 2
E.1 Main Proof of Theorem 2
In this section, we present the proof of Theorem 2, which is obtained by using the RB loss (4) instead of the PL loss (3). Note that this approach is based on the concept of rank breaking (RB), which decomposes (partial) ranking data into individual pairwise comparisons, treats each comparison as independent, and has been extensively studied in previous works [6, 34, 32, 69]. Moreover, this RB approach is applied in the current RLHF for LLM (e.g., Ouyang et al. [59]) and is also studied theoretically in Zhu et al. [96] under the offline setting.
@@ E.1 Main Proof of Theorem 2
$H_{t}:=\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\nabla^{2}\ell_{s}^{(j,k)}(\widehat{\bm{\theta}}_{s}^{(j,k+1)})+\lambda\mathbf{I}_{d},\quad\lambda>0.$
Online confidence bound for RB loss. Now, we introduce the online confidence bound for RB loss. Since the total number of updates up to round $t$ is $\sum_{s=1}^{t}\binom{|S_{s}|}{2}$, a modification of Lemma D.1 yields the following result:
Corollary E.1 (Online confidence bound for RB loss).
@@
Useful definitions. We use the same or similar definitions for the set of warm-up rounds ${\mathcal{T}}^{w}$ (given in Equation (D.1)), the set of large elliptical potential (EP) rounds ${\mathcal{T}}_{0}$ (given in Equation (D.3)), and the regularized covariance matrix $\Lambda_{t}$ (given in Equation (D.2)).
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{\beta_{T+1}(\delta)}\right\},$ (warm-up rounds)
${\mathcal{T}}_{0}$ $:=\left\{t\in[T]:\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}\geq 1.\right\},$ (large EP rounds)
$\Lambda_{t}$ $:=\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}.$
Key Lemmas. We can avoid the $1/\kappa=\mathcal{O}(e^{B})$ dependency in the leading term, thanks to the following lemma.
Lemma E.1.
Let $\Lambda_{t}$ be defined as in Equation (D.2). Set $\lambda=\Omega(d\log(KT/\delta))$. Then, for all $t\in[T]$, with probability at least $1-\delta$, we have
$H_{t}\succeq\frac{1}{10}\Lambda_{t}.$
The proof is deferred to Appendix E.2.1.
Lemma E.2.
Let ${\mathcal{T}}_{0}:=\{t\in[T]:\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}\geq 1.\}$ and ${\mathcal{T}}^{w}=\{t\in[T]:\max_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}\geq\frac{1}{\beta_{T+1}(\delta)}\}$. Define $\kappa:=\frac{e^{-4B}}{4}$. Then, the size of the set ${\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$ is bounded as follows:
$\left|{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}\right|\leq\frac{2}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right).$
The proof is deferred to Appendix E.2.2.
@@
The overall proof structure is similar to that of Theorem 1. We begin with Equation (D.5), but apply Lemma E.2 instead of Lemma D.5. With probability at least $1-\delta$, we have
@@
$\leq\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right)+\frac{8B}{\log(2)T}d\log\left(1+\frac{2K}{\log(2)\lambda}\right)+\frac{8B}{\kappa T}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right)$ (Lemma E.2)
@@
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right).$ (E.1)
@@
To further bound the last term of Equation (E.1), by following the same logic from Equation (D.5) to Equation (D.6), with probability at least $1-\delta$, we obtain
@@
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}$ $\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S^{\star}_{t}|}\sum_{a\in S^{\star}_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$ ($S^{\star}_{t}:=\{\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})\}$)
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}.$ ($S_{t}$ selection rule, Eqn. (9))
To further bound the right-hand side, by applying the Cauchy-Schwartz inequality, we get
$\frac{2\beta_{T+1}(\delta)}{T}$ $\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\frac{1}{|S_{t}|}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}$
@@
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\frac{1}{|S_{t}|}\right)^{2}|S_{t}|}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{H_{t}^{-1}}^{2}}$ (Cauchy-Schwartz ineq.)
$=\frac{2\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\frac{1}{|S_{t}|}\right)^{2}|S_{t}|}\sqrt{10\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}^{2}}$ (Lemma E.1, with probability at least $1-\delta$)
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{10\sum_{t\notin{\mathcal{T}}^{w}}\min\left\{1,\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi\left(x_{t},\bar{a}_{t}\right)\right\|_{\Lambda_{t}^{-1}}^{2}\right\}}$ ($t\notin{\mathcal{T}}_{0}$ and ${\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}\subseteq[T]$)
@@
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{20d\log\left(1+\frac{2K(T-|{\mathcal{T}}^{w}|)}{d\lambda}\right)}$ (Lemma D.3)
@@
$=\mathcal{O}\left(\frac{\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\cdot\sqrt{d\log\left(KT\right)}\right).$ (E.2)
@@
By plugging Equation (E.2) into Equation (E.1) and setting $\beta_{T+1}(\delta)=\mathcal{O}\big(B\sqrt{d\log(KT)}+B\sqrt{\lambda}\big)$, then with probability at least $1-3\delta$, we derive that
@@
Substituting $\delta\leftarrow\frac{\delta}{3}$, we conclude the proof of Theorem 2. ∎
E.2 Proofs of Lemmas for Theorem 2
E.2.1 Proof of Lemma E.1
Proof of Lemma E.1.

### v2

@@ C.2 Proofs of Lemmas for Theorem 1
$\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}$ $\leq\frac{1}{4}\cdot\frac{2}{\lambda}$ ($\dot{\mu}(\cdot)\leq\frac{1}{4}$)
$\leq 1.$ (C.13)
Then, we can derive that
$\sum_{t\in{\mathcal{T}}^{w}}$ $\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$
$=\sum_{t\in{\mathcal{T}}^{w}}\max_{a,a^{\prime}\in\mathcal{A}}\frac{\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)}{\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$
$\leq 4e^{2B}\sum_{t\in{\mathcal{T}}^{w}}\max_{a,a^{\prime}\in\mathcal{A}}\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$ (Assumption 1)
$=4e^{2B}\sum_{t\in{\mathcal{T}}^{w}}\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$
$\leq 8K^{2}e^{6B}\sum_{t\in{\mathcal{T}}^{w}}\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}$ (Eqn. (C.12))
$=8K^{2}e^{6B}\sum_{t\in{\mathcal{T}}^{w}}\min\left\{1,\dot{\mu}\left(\big(\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}\right\}$ (Eqn. (C.13) )
$\leq 16K^{2}e^{6B}d\log\left(1+\frac{2T}{d\lambda}\right).$ (Lemma E.6)
On the other hand, for $t\in{\mathcal{T}}^{w}$, we know that
$\sum_{t\in{\mathcal{T}}^{w}}\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}\geq\frac{|{\mathcal{T}}^{w}|}{18K^{2}\beta_{T+1}(\delta)^{2}}.$
By combining the two results above, and setting $\kappa=e^{-6}$, we obtain
@@
$|{\mathcal{T}}^{w}|\leq\frac{288K^{4}}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2KT}{d\lambda}\right),$
@@
Appendix D Proof of Theorem 2
D.1 Main Proof of Theorem 2
In this section, we present the proof of Theorem 2, which is obtained by using the RB loss (4) instead of the PL loss (3). Note that this approach is based on the concept of rank-breaking (RB), which decomposes (partial) ranking data into individual pairwise comparisons, treats each comparison as independent, and has been extensively studied in previous works [azari2013generalized, khetan2016data, jang2017optimal, saha2024stop]. Moreover, this RB approach is applied in the current RLHF for LLM (e.g., ouyang2022training) and is also studied theoretically in zhu2023principled under the offline setting.
@@ D.1 Main Proof of Theorem 2
$H_{t}$ $:=\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\nabla^{2}\ell_{s}^{(j,k)}(\widehat{\bm{\theta}}_{s}^{(j,k+1)})+\lambda\mathbf{I}_{d}$
$=\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|-1}\!\!\sum_{k=j+1}^{|S_{s}|}\dot{\mu}\left(\big(\phi(x_{t},\sigma_{j})-\phi(x_{t},\sigma_{k})\big)^{\top}\widehat{\bm{\theta}}_{s}^{(j,k+1)}\right)$
$\quad\quad\quad\quad\quad\quad\quad\cdot\big(\phi(x_{t},\sigma_{j})-\phi(x_{t},\sigma_{k})\big)\big(\phi(x_{t},\sigma_{j})-\phi(x_{t},\sigma_{k})\big)^{\top}+\lambda\mathbf{I}_{d}.$
Online confidence bound for RB loss. Now, we introduce the online confidence bound for RB loss. Since the total number of updates up to round $t$ is $\sum_{s=1}^{t}\binom{|S_{s}|}{2}$, a modification of Lemma C.1 yields the following result:
Corollary D.1 (Online confidence bound for RB loss).
@@
Useful definitions. We define the set of warm-up rounds ${\mathcal{T}}^{w}$ in a similar manner to the analysis of Theorem 1:
${\mathcal{T}}^{w}$ $:=\left\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq\frac{1}{\beta_{T+1}(\delta)}\right\}.$ (warm-up rounds)
Moreover, we define the matrix
$\Lambda_{t}=\!\!$ $\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\!\sum_{a,a^{\prime}\in S_{s}}\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{s}\right)\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}$
$+\lambda\mathbf{I}_{d}$ (D.1)
Key lemmas. The relationship between $H_{t}$ and $\Lambda_{t}$ is as follows:
Lemma D.1.
Let $\Lambda_{t}$ be defined as in Equation (D.1). Then, for all $t\in[T]$, we have
$H_{t}\succeq\frac{1}{2e^{2}}\Lambda_{t}.$
The proof is deferred to Appendix D.2.1.
The size of the set ${\mathcal{T}}^{w}$ is bounded as described in the following lemma:
Lemma D.2.
Let ${\mathcal{T}}^{w}=\big\{t\in[T]:\max_{a,a^{\prime}\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}\geq\frac{1}{\beta_{T+1}(\delta)}\big\}$. Define $\kappa:=e^{-6B}$. Set $\lambda\geq 1$. Then, the size of the set ${\mathcal{T}}^{w}$ is bounded as follows:
$\left|{\mathcal{T}}^{w}\right|\leq\frac{32}{\kappa}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{2T}{d\lambda}\right).$
The proof is deferred to Appendix D.2.2. Note that the bound on the cardinality of ${\mathcal{T}}^{w}$ is tighter than in the PL loss case (Lemma C.4), by up to a factor of $K^{4}$ (and up to logarithmic factors in $K$).
@@
The overall proof structure is similar to that of Theorem 1. We start from Equation (C.4), but apply Lemma D.2 instead of Lemma C.4. Then, with probability at least $1-\delta$, we have
@@
$\leq\tilde{\mathcal{O}}\left(\frac{1}{\sqrt{T}}\right)+\mathcal{O}\left(\frac{B}{\kappa T}\beta_{T+1}(\delta)^{2}d\log\left(1+\frac{T}{d\lambda}\right)\right)$ (Lemma D.2)
@@
$+\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right).$ (D.2)
@@
To further bound the last term of Equation (D.2), by following the similar reasoning from Equation (C.4) to Equation (C.6), with probability at least $1-\delta$, we obtain
$\sum_{t\notin{\mathcal{T}}^{w}}$ $\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
@@
$\leq e(1+e)^{2}\sum_{t\notin{\mathcal{T}}^{w}}\frac{1}{|S^{\star}_{t}|}\sum_{a,a^{\prime}\in S^{\star}_{t}}\dot{\mu}\big(\psi_{t}(a,a^{\prime})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,a^{\prime})\right\|_{H_{t}^{-1}}$ ($S^{\star}_{t}:=\{\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})\}$, and let $\psi_{t}(a,a^{\prime})=\phi(x_{t},a)-\phi(x_{t},a^{\prime})$)
@@
$\leq e(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{|S_{t}|}{|S^{\star}_{t}|}\sum_{a,a^{\prime}\in S^{\star}_{t}}\dot{\mu}\big(\psi_{t}(a,a^{\prime})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,a^{\prime})\right\|_{H_{t}^{-1}}^{2}}$ (Cauchy-Schwartz ineq., and ${\mathcal{T}}^{w}\subseteq[T]$)
$\leq e(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{\cancel{|S_{t}|}}{\cancel{|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}}|}}\sum_{a,a^{\prime}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}}}\dot{\mu}\big(\psi_{t}(a,a^{\prime})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,a^{\prime})\right\|_{H_{t}^{-1}}^{2}}$
@@
$\leq\sqrt{2}e^{2}(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\sum_{a,a^{\prime}\in S_{t}}\dot{\mu}\big(\psi_{t}(a,a^{\prime})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,a^{\prime})\right\|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Lambda_{t}^{-1}}}^{2}}.$ (Lemma D.1)
Moreover, by setting $\lambda=d\log T$ and $T\geq e^{K^{2}/d}$, for any $t$, we have
$\sum_{a,a^{\prime}\in S_{t}}\dot{\mu}\big(\psi_{t}(a,a^{\prime})^{\top}\widehat{\bm{\theta}}_{t}\big)\left\|\psi_{t}(a,a^{\prime})\right\|_{\Lambda_{t}^{-1}}^{2}\leq\frac{2K^{2}}{\lambda}\leq 2.$ (D.3)
Hence, by the elliptical potential lemma (Lemma E.6), we derive that
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}$ $\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
@@
By plugging Equation (D.4) into Equation (D.2) and setting $\beta_{T+1}(\delta)=\mathcal{O}\big(B\sqrt{d\log(KT)}+B\sqrt{\lambda}\big)$, then, with probability at least $1-2\delta$, we derive that
@@
Substituting $\delta\leftarrow\frac{\delta}{2}$, we conclude the proof of Theorem 2. ∎
D.2 Proofs of Lemmas for Theorem 2
D.2.1 Proof of Lemma D.1
Proof of Lemma D.1.